# Denominators in Lusztig’s asymptotic Hecke algebra via the Plancherel formula

### v5

@@ Abstract
Let $\tilde{W}$ be an extended affine Weyl group, $\mathbf{H}$ be the corresponding affine Hecke algebra over the ring $\mathbb{C}[\mathbf{q}^{\frac{1}{2}},\mathbf{q}^{-\frac{1}{2}}]$, and $J$ be Lusztig’s asymptotic Hecke algebra, viewed as a based ring with basis $\{t_{w}\}$. Viewing $J$ as a subalgebra of the $(\mathbf{q}^{-\frac{1}{2}})$-adic completion of $\mathbf{H}$ via Lusztig’s map $\phi$, we use Harish-Chandra’s Plancherel formula for $p$-adic groups to show that the coefficient of $T_{x}$ in $t_{w}$ is a rational function of $\mathbf{q}$, with denominator depending only on the two-sided cell containing $w$, and dividing a power of the Poincaré polynomial of the finite Weyl group. As an application, we conjecture that these denominators encode more detailed information about the failure of the Kazhdan-Lusztig classification at roots of the Poincaré polynomial than is currently known.
@@ 1 Introduction
Let $\tilde{W}$ be an affine Weyl group or extended affine Weyl group, and let $\mathbf{H}$ be its associated Hecke algebra over $\mathcal{A}:=\mathbb{C}[\mathbf{q}^{1/2},\mathbf{q}^{-1/2}]$, where $\mathbf{q}$ is a formal variable. The representation theory of $\mathbf{H}$ is very well understood, behaving well and uniformly when $\mathbf{q}$ is specialized to any $q\in\mathbb{C}^{\times}$ that is not a root of the Poincaré polynomial $P_{W}$ of the finite Weyl group $W\subset\tilde{W}$.
@@
One way Lusztig expressed the uniformity in $q$ of the representation theory of the various algebras $\mathbf{H}|_{\mathbf{q}=q}$ is via the asymptotic Hecke algebra $J$. This is a $\mathbb{C}$-algebra (in fact, a $\mathbb{Z}$-algebra) $J$ with distinguished basis $\{t_{w}\}_{w\in\tilde{W}}$, equipped with an injection $\phi\colon\mathbf{H}\mathrel{\hookrightarrow}J\otimes_{\mathbb{C}}\mathcal{A}$. In this way there is a map from $J$-modules to $\mathbf{H}$-modules, and Lusztig has shown in [affineIII] and [affineIV] that when $q$ is not a root of unity (other than $1$), that the specialized map $\phi_{q}$ induces a bijection between simple $\mathbf{H}|_{\mathbf{q}=q}$-modules and simple $J$-modules, these last being defined over $\mathbb{C}$. Moreover, he showed that when $P_{W}(q)\neq 0$, the map $\phi_{q}$ induces an isomorphism
@@
$()^{\dagger}\circ\phi^{-1}(t_{w})=\sum_{x\in\tilde{W}}a_{x,w}T_{x},$ (1)
where each $a_{x,w}$ is a formal Laurent series in $\mathbf{q}^{-1/2}$, and $()^{\dagger}$ is the Goldman involution.
@@ 1.1 The asymptotic Hecke algebra and $p$-adic groups
More precisely, let $\mathbf{G}$ be a connected reductive group defined and split over a non-archimedean local field $F$ whose extended affine Weyl group is $\tilde{W}$. Then in [BK], Braverman and Kazhdan define a map expressing $J$ as sitting between the Iwahori-Hecke algebra of $G=\mathbf{G}(F)$ and the Harish-Chandra Schwartz algebra $\mathcal{C}$, and propose a spectral characterization of $J$ via the operator Payley-Wiener theorem, obtaining a diagram
@@ 1.1.1 Denominators in the affine Hecke algebra and injectivity of $\eta$
[…]
@@
Let $\tilde{W}$ be an affine Weyl group, $\mathbf{H}$ its affine Hecke algebra over $\mathcal{A}$, and $J$ its asymptotic Hecke algebra. Let $\phi\colon\mathbf{H}\mathrel{\hookrightarrow}J\otimes_{\mathbb{C}}\mathcal{A}$ be Lusztig’s map.
@@
    For all $x,w\in\tilde{W}$, $a_{x,w}$ is a rational function of $\mathbf{q}$. The denominator of $a_{x,w}$ is independent of $x$. As a function of $w$, it is constant on two-sided cells.
@@
    There exists $N_{\tilde{W}}\in\mathbb{N}$ such that upon writing
    $()^{\dagger}\circ\phi^{-1}(t_{w})=\sum_{x\in\tilde{W}}a_{x,w}T_{x},$
@@
    $P_{W}(\mathbf{q})^{N_{\tilde{W}}}a_{x,w}\in\mathcal{A}$
    for all $x,w\in\tilde{W}$.
@@
In [D], the author proved Theorem 2 in type $\tilde{A}_{1}$, but with different conventions. To translate to the conventions of this paper, the reader should replace $j$ with the Goldman involution $()^{\dagger}$, and the completion of $\mathbf{H}$ with respect to the $C_{w}$ basis and positive powers of $\mathbf{q}^{1/2}$ with the completion of $\mathbf{H}$ with respect to the basis $\{(-1)^{\ell(w)}C^{\prime}_{w}\}_{w\in\tilde{W}}$ and negative powers of $\mathbf{q}^{1/2}$. Note also that we write $a_{x,w}$ instead of $a_{w,x}$ as in [D]. In [N], Neunhöffer described the coefficients $a_{x,w}$ for finite Weyl groups.
In future, it would be desirable to also treat the case of unequal parameters, where a result like that of [BDD] governing denominators is not yet available. On the other hand, when $G=\mathrm{GL}_{n}$, we are able to be slightly more precise than Theorem 2, while also not appealing to [BDD]. For this reason we treat the case $G=\mathrm{GL}_{n}$ separately as Theorem 13.
@@ 1.1.2 Denominators in the affine Hecke and the Kazhdan-Lusztig classification at roots of unity
By Theorem 2, the coefficients $a_{1,d}$ have poles at every root of $P_{W}$, for all distinguished involutions $d$ in the lowest two-sided cell $\mathbf{c}_{0}$. On the other hand, as we show in Example 7, there do exist cells $\mathbf{c}\neq\mathbf{c}_{0}$ such that the coefficients $a_{x,w}$ are nonsingular at certain roots of $P_{W}$, for all $w\in\mathbf{c}$ and $x\in\tilde{W}$. We encode the hope that this is no accident as
@@
Let $\tilde{W}$ be an affine Weyl group, and let $q\in\mathbb{C}^{\times}$ be a root of $P_{W}$. Let $\mathbf{c}$ be a two-sided cell such that if $w\in\mathbf{c}$, then $a_{x,w}$ does not have a pole at $\mathbf{q}=q$ for any $x\in\tilde{W}$. Let $u=u(\mathbf{c})$. Then every standard module $K(u,s,\rho,q)$ has a unique simple quotient $L=L(u,s,\rho,q)$ such that $a(L)=a(E)$, where $E$ is the simple $J$ module corresponding to $K(u,s,\rho,q)$, in the notation of [affineIV]. Two such simple modules are isomorphic if and only if their corresponding triples are conjugate.
@@ 1.1.3 Application: representations with parahoric-fixed vectors
In Section 4, we use the existence of the action of the asymptotic Hecke algebra on tempered $G$-representations to give a simple criterion for the existence of vectors fixed under a parahoric subgroup of $G$: thm existence of parahoric-fixed vectors
@@
Let $\pi=K(u,s,\rho)$ be a simple tempered $I$-spherical representation of $G$. Let $\mathcal{P}$ be a parahoric subgroup of $G$ and let $w_{\mathcal{P}}$ be the longest element in the corresponding subgroup of $\tilde{W}$. Let $\mathcal{B}_{u}^{\vee}$ be the Springer fibre for $u$.
@@
By [affineIV, Theorem 4.8(d)] and the proof of [affineIV, Lemma 5.5], every two-sided cell contains a distinguished involution contained in a finite parabolic subgroup of $\tilde{W}$, but not every distinguished involution of a finite Coxeter group is the longest word of a parabolic subgroup, i.e. is of the form $w_{\mathcal{P}}$; approximately half of two-sided cells of the finite Weyl group $W\subset\tilde{W}$ do not contain any distinguished involutions contained in proper parabolic subgroups, because of the cell-preserving bijection $w\mapsto w_{0}w$. For example, this happens for the second-lowest cell for $E_{8}$.
@@ 1.3 The affine Hecke algebra
Let $\mathbf{G}$ be a connected reductive algebraic group defined and split over $F$, $\mathbf{A}$ a maximal torus of $\mathbf{G}$, and $X_{*}=X_{*}(\mathbf{A})$ the cocharacter lattice of $\mathbf{A}$. Let $\mathbf{N}$ be unipotent radical of a chosen Borel subgroup $\mathbf{B}$, so that $\mathbf{B}=\mathbf{A}\mathbf{N}$. Let $W$ be the finite Weyl group of $\mathbf{G}$, and $\tilde{W}=W\ltimes X_{*}(\mathbf{A})$ be the extended affine Weyl group. Write $S$ for the set of simple reflections in $\tilde{W}$. Let $\mathbf{G}^{\vee}$ be the Langlands dual group of $\mathbf{G}$, taken over $\mathbb{C}$. We write $G=\mathbf{G}(F)$, $A=\mathbf{A}(F)$, etc. Where there is no danger of confusion, we also write $G^{\vee}$ for $\mathbf{G}^{\vee}(\mathbb{C})$, $M^{\vee}$ for $\mathbf{M}^{\vee}(\mathbb{C})$, etc. Let $K$ be the maximal compact subgroup $\mathbf{G}(\mathcal{O})$. Also let $I$ be the Iwahori subgroup of $G$ that is the preimage of $\mathbf{B}(\mathbb{F}_{q})$. We sometimes write $P_{\mathbf{G}/\mathbf{B}}$ for $P_{W}$, as this polynomial is also the Poincaré polynomial of the flag variety $(\mathbf{G}/\mathbf{B})(\mathbb{C})$.
We write $\mathbf{H}$ for the affine Hecke algebra of $\tilde{W}$. It is a unital associative algebra over the ring $\mathcal{A}=\mathbb{C}[\mathbf{q}^{\frac{1}{2}},\mathbf{q}^{-\frac{1}{2}}]$ (in fact, it is defined over $\mathbb{Z}[\mathbf{q}^{\frac{1}{2}},\mathbf{q}^{-\frac{1}{2}}]$ but we will work over $\mathbb{C}$ to avoid having to introduce extra notation later), where $\mathbf{q}^{\frac{1}{2}}$ is a formal variable. We will think of $\mathbb{C}^{\times}$ as $\operatorname{Spec}\mathcal{A}$. The algebra $\mathbf{H}$ has the Coxeter presentation with standard basis $\{T_{w}\}_{w\in\tilde{W}}$ with $T_{w}T_{w^{\prime}}=T_{ww^{\prime}}$ if $\ell(ww^{\prime})=\ell(w)+\ell(w^{\prime})$ and quadratic relation $(T_{s}+1)(T_{s}-\mathbf{q})=0$ for $s\in S$. We write $\theta_{\lambda}$ for the generators of the Bernstein subalgebra.
Recall from [KL79] the two Kazhdan-Lusztig bases $\{C_{w}\}_{w\in\tilde{W}}$ and $\{C^{\prime}_{w}\}_{w\in\tilde{W}}$ of $\mathbf{H}$ and write $C_{x}C_{y}=\sum_{z\in\tilde{W}}h_{x,y,z}C_{z}$. The inverse Kazhdan-Lusztig polynomials $Q_{y,x}$ are the unique family of polynomials satisfying

### v6

@@ Abstract
Let $W_{\mathrm{aff}}$ be an extended affine Weyl group, $\mathbf{H}$ be the corresponding affine Hecke algebra over the ring $\mathbb{C}[\mathbf{q}^{\frac{1}{2}},\mathbf{q}^{-\frac{1}{2}}]$, and $J$ be Lusztig’s asymptotic Hecke algebra, viewed as a based ring with basis $\{t_{w}\}$. Viewing $J$ as a subalgebra of the $(\mathbf{q}^{-\frac{1}{2}})$-adic completion of $\mathbf{H}$ via Lusztig’s map $\phi$, we use Harish-Chandra’s Plancherel formula for $p$-adic groups to show that the coefficient of $T_{x}$ in $t_{w}$ is a rational function of $\mathbf{q}$, with denominator depending only on the two-sided cell containing $w$, and dividing a power of the Poincaré polynomial of the finite Weyl group. As an application, we conjecture that these denominators encode more detailed information about the failure of the Kazhdan-Lusztig classification at roots of the Poincaré polynomial than is currently known.
@@ 1 Introduction
Let $W_{\mathrm{aff}}$ be an affine Weyl group or extended affine Weyl group, and let $\mathbf{H}$ be its associated Hecke algebra over $\mathcal{A}:=\mathbb{C}[\mathbf{q}^{1/2},\mathbf{q}^{-1/2}]$, where $\mathbf{q}$ is a formal variable. The representation theory of $\mathbf{H}$ is very well understood, behaving well and uniformly when $\mathbf{q}$ is specialized to any $q\in\mathbb{C}^{\times}$ that is not a root of the Poincaré polynomial $P_{W}$ of the finite Weyl group $W\subset W_{\mathrm{aff}}$.
@@
One way Lusztig expressed the uniformity in $q$ of the representation theory of the various algebras $\mathbf{H}|_{\mathbf{q}=q}$ is via the asymptotic Hecke algebra $J$. This is a $\mathbb{C}$-algebra (in fact, a $\mathbb{Z}$-algebra) $J$ with distinguished basis $\{t_{w}\}_{w\in W_{\mathrm{aff}}}$, equipped with an injection $\phi\colon\mathbf{H}\mathrel{\hookrightarrow}J\otimes_{\mathbb{C}}\mathcal{A}$. In this way there is a map from $J$-modules to $\mathbf{H}$-modules, and Lusztig has shown in [affineIII] and [affineIV] that when $q$ is not a root of unity (other than $1$), that the specialized map $\phi_{q}$ induces a bijection between simple $\mathbf{H}|_{\mathbf{q}=q}$-modules and simple $J$-modules, these last being defined over $\mathbb{C}$. Moreover, he showed that when $P_{W}(q)\neq 0$, the map $\phi_{q}$ induces an isomorphism
@@
$^{\dagger}(-)\circ\phi^{-1}(t_{w})=\sum_{x\in W_{\mathrm{aff}}}a_{x,w}T_{x},$ (1)
where each $a_{x,w}$ is a formal Laurent series in $\mathbf{q}^{-1/2}$, and $^{\dagger}(-)$ is the involution of $\mathbf{H}$ defined in Definition 6. It agrees with the Goldman involution of $\mathbf{H}$ when $G$ is simply-connected. In this paper we will almost exclusively work with $\phi\circ{}^{\dagger}(-)$, for reasons explained in Section 1.1.
@@ 1.1 The asymptotic Hecke algebra and $p$-adic groups
More precisely, let $\mathbf{G}$ be a connected reductive group defined and split over a non-archimedean local field $F$ whose extended affine Weyl group is $W_{\mathrm{aff}}$. Then in [BK], Braverman and Kazhdan define a map expressing $J$ as sitting between the Iwahori-Hecke algebra of $G=\mathbf{G}(F)$ and the Harish-Chandra Schwartz algebra $\mathcal{C}$, and propose a spectral characterization of $J$ via the operator Payley-Wiener theorem, obtaining a diagram
@@ 1.1.1 Denominators in the affine Hecke algebra and injectivity of $\eta$
In the first part of this paper we prove that $\eta$ is injective. Along the way, we prove in Proposition 3 that $\tilde{\phi}$ is essentially the map $\phi^{-1}$. Our strategy is to determine that the Schwartz functions $f_{w}$ on the $p$-adic group $G$ satisfy the statements of Theorem 2 below, and are in addition sufficiently well-behaved so as to lift to elements of a certain completion $\mathbfcal{H}^{-}$ of $\mathbf{H}$, thus defining a map
$\phi_{1}\colon J\otimes_{\mathbb{C}}\mathcal{A}\to\mathbfcal{H}^{-}.$
We therefore obtain two maps $J\subset J\otimes_{\mathbb{C}}\mathcal{A}\to\mathbfcal{H}^{-}$: the inverse $^{\dagger}(-)\circ\phi^{-1}$ of Lusztig’s map, and our map $\phi_{1}$ induced by the construction in [BK]. We prove in Proposition 3 that these maps agree, at which point Theorem 13 and the first statement of Theorem 2 follow.
@@
Let $W_{\mathrm{aff}}$ be an affine Weyl group, $\mathbf{H}$ its affine Hecke algebra over $\mathcal{A}$, and $J$ its asymptotic Hecke algebra. Let $\phi\colon\mathbf{H}\mathrel{\hookrightarrow}J\otimes_{\mathbb{C}}\mathcal{A}$ be Lusztig’s map.
@@
    For all $x,w\in W_{\mathrm{aff}}$, $a_{x,w}$ is a rational function of $\mathbf{q}$. The denominator of $a_{x,w}$ is independent of $x$. As a function of $w$, it is constant on two-sided cells.
@@
    There exists $N_{W_{\mathrm{aff}}}\in\mathbb{N}$ such that upon writing
    $^{\dagger}(-)\circ\phi^{-1}(t_{w})=\sum_{x\in W_{\mathrm{aff}}}a_{x,w}T_{x},$
@@
    $P_{W}(\mathbf{q})^{N_{W_{\mathrm{aff}}}}a_{x,w}\in\mathcal{A}$
    for all $x,w\in W_{\mathrm{aff}}$.
@@
In [D], the author proved Theorem 2 in type $\tilde{A}_{1}$, but with different conventions. To translate to the conventions of this paper, the reader should replace $j$ with the involution $^{\dagger}(-)$, and the completion of $\mathbf{H}$ with respect to the $C_{w}$ basis and positive powers of $\mathbf{q}^{1/2}$ with the completion of $\mathbf{H}$ with respect to the basis $\{(-1)^{\ell(w)}C^{\prime}_{w}\}_{w\in W_{\mathrm{aff}}}$ and negative powers of $\mathbf{q}^{1/2}$. Note also that we write $a_{x,w}$ instead of $a_{w,x}$ as in [D]. In [N], Neunhöffer described the coefficients $a_{x,w}$ for finite Weyl groups.
In future, it would be desirable to also treat the case of unequal parameters, where a result like that of [BDD] governing denominators is not yet available. On the other hand, when $G=\mathrm{GL}_{n}$, we are able to be slightly more precise than Theorem 2, while also not appealing to [BDD]. For this reason we treat the case $G=\mathrm{GL}_{n}$ separately as Theorem 13 in Section 3.4.
@@ 1.1.2 Denominators in the affine Hecke and the Kazhdan-Lusztig classification at roots of unity
By Theorem 2, the coefficients $a_{1,d}$ have poles at every root of $P_{W}$, for all distinguished involutions $d$ in the lowest two-sided cell $\mathbf{c}_{0}$. On the other hand, as we show in Example 7, there do exist cells $\mathbf{c}\neq\mathbf{c}_{0}$ such that the coefficients $a_{x,w}$ are nonsingular at certain roots of $P_{W}$, for all $w\in\mathbf{c}$ and $x\in W_{\mathrm{aff}}$. We encode the hope that this is no accident as
@@
Let $\tilde{W}$ be an affine Weyl group, and let $q\in\mathbb{C}^{\times}$ be a root of $P_{W}$. Let $\mathbf{c}$ be a two-sided cell such that if $w\in\mathbf{c}$, then $a_{x,w}$ does not have a pole at $\mathbf{q}=q$ for any $x\in W_{\mathrm{aff}}$. Let $u=u(\mathbf{c})$. Let $K(u,s,\rho)$ be a standard module in the notation of [KLDeligneLanglands]. Then the module $^{\dagger}{}^{*}K(u,s,\rho,q)$ (see Definition 6 (a), (b) and the discussion following Theorem 7) has a unique simple quotient $L=L(u,s,\rho,q)$ such that $a(L)=a(E)$, where $E$ is the simple $J$ module corresponding via $\phi$ to $(u,s,\rho)$ under [affineIV, Thm. 4.2]. Two such simple modules are isomorphic if and only if their corresponding triples are conjugate.
@@ 1.1.3 Application: representations with parahoric-fixed vectors
In Section 4, we use the existence of the action of the asymptotic Hecke algebra on tempered $G$-representations to give a simple criterion for the existence of vectors fixed under a parahoric subgroup of $G$:
@@
Let $\pi=K(u,s,\rho)$ be a simple tempered $I$-spherical representation of $G$. Let $\mathcal{P}$ be a parahoric subgroup of $G$ and let $w_{\mathcal{P}}$ be the longest element in the corresponding subgroup of $W_{\mathrm{aff}}$. Let $\mathcal{B}_{u}^{\vee}$ be the Springer fibre for $u$.
@@
By [affineIV, Theorem 4.8(d)] and the proof of [affineIV, Lemma 5.5], every two-sided cell contains a distinguished involution contained in a finite parabolic subgroup of $W_{\mathrm{aff}}$, but not every distinguished involution of a finite Coxeter group is the longest word of a parabolic subgroup, i.e. is of the form $w_{\mathcal{P}}$; approximately half of two-sided cells of the finite Weyl group $W\subset W_{\mathrm{aff}}$ do not contain any distinguished involutions contained in proper parabolic subgroups, because of the cell-preserving bijection $w\mapsto w_{0}w$. For example, this happens for the second-lowest cell for $E_{8}$.
@@ 1.3 The affine Hecke algebra
Let $\mathbf{G}$ be a connected reductive algebraic group defined and split over $F$, $\mathbf{A}$ a maximal $F$-split torus of $\mathbf{G}$, and $X_{*}=X_{*}(\mathbf{A})$ the cocharacter lattice of $\mathbf{A}$. Let $\pi_{1}(G)=X_{*}/\mathbb{Z}\Phi^{\vee}$ be the fundamental group, the quotient of the cocharacter lattice by the coroot lattice. Let $\mathbf{N}$ be unipotent radical of a chosen Borel subgroup $\mathbf{B}$, so that $\mathbf{B}=\mathbf{A}\mathbf{N}$. Let $W$ be the finite Weyl group of $\mathbf{G}$, and $W_{\mathrm{aff}}=W\ltimes X_{*}(\mathbf{A})$ be the extended affine Weyl group. Write $S$ for the set of simple reflections in $W_{\mathrm{aff}}$. Let $\mathbf{G}^{\vee}$ be the Langlands dual group of $\mathbf{G}$, taken over $\mathbb{C}$. We write $G=\mathbf{G}(F)$, $A=\mathbf{A}(F)$, etc. Where there is no danger of confusion, we also write $G^{\vee}$ for $\mathbf{G}^{\vee}(\mathbb{C})$, $M^{\vee}$ for $\mathbf{M}^{\vee}(\mathbb{C})$, etc. Let $K$ be the maximal compact subgroup $\mathbf{G}(\mathcal{O})$. Also let $I$ be the Iwahori subgroup of $G$ that is the preimage of $\mathbf{B}(\mathbb{F}_{q})$ in $K$. We sometimes write $P_{\mathbf{G}/\mathbf{B}}$ for $P_{W}$, as this polynomial is also the Poincaré polynomial of the flag variety $(\mathbf{G}/\mathbf{B})(\mathbb{C})$.
We write $\mathbf{H}$ for the affine Hecke algebra of $W_{\mathrm{aff}}$. It is a unital associative algebra over the ring $\mathcal{A}=\mathbb{C}[\mathbf{q}^{\frac{1}{2}},\mathbf{q}^{-\frac{1}{2}}]$ (in fact, it is defined over $\mathbb{Z}[\mathbf{q}^{\frac{1}{2}},\mathbf{q}^{-\frac{1}{2}}]$ but we will work over $\mathbb{C}$ to avoid having to introduce extra notation later), where $\mathbf{q}^{\frac{1}{2}}$ is a formal variable. We will think of $\mathbb{C}^{\times}$ as $\operatorname{Spec}\mathcal{A}$. The algebra $\mathbf{H}$ has the Coxeter presentation with standard basis $\{T_{w}\}_{w\in\tilde{W}}$ with $T_{w}T_{w^{\prime}}=T_{ww^{\prime}}$ if $\ell(ww^{\prime})=\ell(w)+\ell(w^{\prime})$ and quadratic relation $(T_{s}+1)(T_{s}-\mathbf{q})=0$ for $s\in S$. We write $\theta_{\lambda}$ for the generators of the Bernstein subalgebra.
Recall from [KL79] the two Kazhdan-Lusztig bases $\{C_{w}\}_{w\in W_{\mathrm{aff}}}$ and $\{C^{\prime}_{w}\}_{w\in W_{\mathrm{aff}}}$ of $\mathbf{H}$, where
$C^{\prime}_{w}=\mathbf{q}^{-\frac{\ell(w)}{2}}\sum_{x\leq w}P_{x,w}(\mathbf{q})T_{x}$
for the Kazhdan-Lusztig polynomials $P_{x,w}$. Write $C_{x}C_{y}=\sum_{z\in W_{\mathrm{aff}}}h_{x,y,z}C_{z}$. The inverse Kazhdan-Lusztig polynomials $Q_{y,x}$ are the unique family of polynomials satisfying
@@
If
$\varphi\colon(W_{\mathrm{aff}},S)\to(W_{\mathrm{aff}},S)$
is a Coxeter group automorphism of $W_{\mathrm{aff}}$, then
$T_{w}\mapsto T_{\varphi(w)}$
is an algebra automorphism of $\mathbf{H}$ commuting with the bar involution, and therefore given equivalently by
$C^{\prime}_{w}\mapsto C^{\prime}_{\varphi(w)}$
and
$C_{w}\mapsto C_{\varphi(w)}.$ (3)